Kutztown University of Pennsylvania Department of Mathematics

Colloquium

3:30 p.m.

March 17, 2009

Lytle Hall 228

Mathematical Cranks

Dr. Lyn Phy

Associate Professor of Mathematics Department of Mathematics Kutztown University of Pennsylvania

Abstract

A ”mathematical crank” is a person who thinks she has done something impossible, like proving Euclid’s Parallel Postulate; or thinks she has done something she has not, like finding a simple proof for the Four Color Theorem; or has eccentric views about the power of mathematics, like believing that the trisection of an angle would enable gold to be made out of lead. In this light-hearted talk, we will explore several such cranks with the hopes that contemplating the odd, peculiar, and bizarre can be enlightening as well as entertaining.

Refreshments served between 3:00 and 3:30 p.m.

The talk begins at 3:30 p.m.

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache III, Chair